OpenStudy (lexixd):

Solve for x. x - 3/8 = - 5/8

iYuko (iyuko):

\[\frac{ x-3 }{ 8 }=-\frac{ 5 }{ 8 }\]

rebeccaxhawaii (rebeccaxhawaii):

so use the reverse effect -5/8+3/8 and that will give you your answer